# Conditions for enhancement of chemical reactions in gas phase inside a dark cavity

### v1

@@ Abstract
Enhancing chemical reactions, such as $A+B\to[\textit{activated complex}]^{\#}\to C+D$, in gas phase through its coupling to quantum-electrodynamics (QED) modes in a dark cavity is investigated. The main result is that the enhancement of the reaction rate by a dark cavity is for asymmetric reactions (products different from reactants.) Notice that in addition to the cavity been dark, the reactants are in their ground electronic and vibrational states, i.e., it is indeed dark. Theoretical derivation, utilizing the non-Hermitian formalism of quantum mechanics (NHQM), provides conditions and guidelines for selecting the proper type of reactions that can be enhanced by a dark cavity. Nevertheless, the time-dependent simulations of such experiments can be carried out using the standard (Hermitian) scattering theory (but including the conditions derived via NHQM). We believe that this work opens a gate to new types of studies and hopefully helps to close the gap between theory and experiments in this fascinating, relatively new field of research. As an example, we demonstrate that the asymmetric reaction rates of $O+D_{2}\to[ODD]^{\#}\to OD+D$ and $H+ArCl\to[ArHCl]^{\#}\to H+Ar+Cl$ can be enhanced by a dark cavity. Contrary, the dark cavity effect on the symmetric reaction of hydrogen exchange in methane will be negligible.
@@ I Motivation
The ability to control, i.e., to slow down or enhance chemical reactions, by a seemingly simple setup of reactions inside a cavity made of two parallel mirrors is fascinating. Unfortunately, currently, theory and experiment have not yet fully converged. See for example Ref.5 where they failed to reproduce the enhanced rate of cyanate ion hydrolysis reported by Hiura et al[6]. This might be due to the recent findings indicate that the cavity can introduce friction between the molecules and the solvent. This friction can enhance the reaction rate when it is low and inhibit it when it becomes significant. See for example study of the cis-trans isomerization of HONO by Sun and Vendrell[7] and references therein.
@@
The theory presented below is based on the non-Hermitian formalism of quantum mechanics (NHQM), which has previously been utilized in studying the cavity effects in chemistry. It led to the discovery of novel phenomena, such as the novel mechanism for generating high-purity single-photon emission at high repetition rates from a coherent light source[8], and the conditions enabling the upper bound polariton to penetrate into the continuum and become finite lifetime metastable state[9]. NHQM enables the calculations of the complex discrete poles of the scattering matrix, i.e., a single NHQM eigenstate instead of a Hermitian wavepacket. We will use these poles as a basis set in calculating cross sections and reaction rates. The advantage of this non-standard approach is in the ability to isolate the structures in the cross sections from the continuum background. A single time independent resonance complex pole may include all measurable information of a specific quantity in spite of the fact that in hermitian quantum mechanics one needs to carryout wave packet dynamical calculations to describe this phenomena. This theoretical approach directs us toward identifying which kinds of collision reactions in gas phase can be enhanced by conducting them between two mirrors where the distance between them is carefully chosen. Recently we have shown that by using the complex poles as a basis set a remarkable agreement between theory and experiment was obtained. This was shown without using any fitting parameters, where the only input parameters for the cold molecular collision between triplet excited Helium and hydrogen molecule where the charges and masses of electrons and nuclei and the Planck constant[10, 11]. We expect to get similar high agreement between theory and experiment in our study of collision reactions such as $A+BC\to AB+C$ where a finite lifetime activated complex $[A...B...C]^{\#}$ is formed and its life time is associated with the complex resonance poles of the scattering matrix. Notice that these complex poles describe predissociation of metastable states of the activated complex. This is to differ from the isomerization reactions mentioned above[7] where the systems is a bound system and the poles do not appear in the complex energy plane but are on the real energy axis. This is an essential point for the theory presented below that provides the conditions for which collision reactions in gas phase (no frictions) can be enhanced in dark cavity due to the QED states in the vacuum.
@@
First, the principles of the mechanism that enable the enhancement of reactions in the gas phase by QED waves in a cavity are described intuitively without equations using the stadard (hermitian) quantum mechanics formulation. Second, a brief overview of the fundamental concepts pertinent to this study will be provided. This will encompass the definition of discrete complex poles within the scattering matrix and its role as a basis set. In addition, it describes the predissociation-resonance metastale states of the activated complex $[ABC]^{\#}$, and the nuclear adiabatic theorem that simplified the calculations of the finite-lifetime resonances of $[ABC]^{\#}$. Second, the conditions that should be satisfied in order to enhance the rate of a reaction in a cavity will be derived. We demonstrate that the reaction rate for $O+D_{2}\to[ODD]^{\#}\to OD+D$ can be increased by placing it inside a dark cavity. The conditions for the enhancement are obtained through tuning the distances between the two cavity-mirrors. This tuning supports a QED wave that couples two resonance states, the occupied state, referred to a transition state (TS) and an unpopulated one referred to a dynamical barrier (DB) resonance state. The DB state remains unpopulated when the distance between the two mirrors is detuned. Third, we will discuss a possible feasible experiments based on the conditions derived herein. Last we conclude.
@@ II An intuitive (hermitian) explanation of the cavity-induced reaction rate enhancement
Lets assume we carryout standard wavepacket calculations to describe the transition from reactants to products on a single electronic multi-dimensional potential energy surface (PES). Although in this paper we employ the non-Hermitian formalism, in which this dynamics is described differently using the non-Hermitian time-independent Hamiltonian. In this Section we do not provide a non-Hermitian explanation since the conclusions are the same. In order to cross the potential barrier that separates the reactants from the products, the collision energy must be at least equal to the energy of the barrier. This is because the transition effect due to tunneling is negligible. However, when the vibrational frequency of the colliding molecules changes substantially during the collision then the transition from reactants to products can be made at a lower collision energy. This is enabled due to the transfer of energy from the transnational motion along the reaction coordinate to excite a bound vibrational state, which is perpendicular to the reaction coordinate. This description is equivalent to bypassing the static potential barrier along the 1D reaction coordinate but in the multidimensional potential energy surface (PES). In the multidimensional PES dynamics, the transition from reactants to products is through a saddle point, which is the transition state. When the reactants are far apart the system is described as a delocalized wavepacket, it is a continuum function along the reaction coordinate. However, around the PES saddle-point the wavepacket becomes a Gaussian like function (without a node along the coordinates, this important for identifying the transition state). Notice that in the perpendicular direction the transition state is in a highly excited vibrational state.
Thus, the reaction rate depends on the strength of the coupling between the transnational motion along the reaction coordinate to the bound vibrations perpendicular to the reaction coordinate. Above we described how the strength of this coupling is due to the changes in the frequency of the vibration along the reaction coordinate. Inside the cavity the coupling of the molecular modes with the QED modes enhance the strength of this coupling and thereby enahnce the rate of transition from reactants to products. Within NHQM it is possible to determine a suitable reaction for accelerating the reaction rate in a dark cavity without performing simulations of the experiment. In addition, it allows determination of the optimal distance between the mirrors in the cavity that will enable the optimal coupling described above.
@@ III Brief overview of the fundamental concepts pertinent to this study
Let us focus here on a collision of atom $A$ with diatom $BC$ (reactants) to form a finite lifetime activated complex $[ABC]^{\#}$, and then proceeds to yield the products $AB+C$. Reaction rates[12] and its cumulative reaction probabilities[13] can be calculated by the standard (Hermitian) quantum scattering theory. In particular, Ref.13 presents the transition from Lippmann-Schwinger formula for transition probability to Flux operator, calculated at any given divided surface. In standard quantum scattering theory the Green operator is defined as $\hat{G}^{+}(E)=\lim_{\epsilon\to 0}(E+i\epsilon-\hat{H}_{ABC})^{-1}$ where $\hat{H}_{ABC}$ is the real physical Hamiltonian for the $A+BC\to AB+C$ reaction. Notice that the term $i\epsilon$ in the Green operator introduces outgoing wave boundary condition into $|\Phi\rangle=\hat{G}^{+}(E)|\phi\rangle$ (see solution to exercise 10.1 in Ref.14). Thus, the green operator is a non-Hermitian operator. Let us quote from Ref.15: “The branching of quantum mechanics to standard (Hermitian) formalism and non-Hermitian (NH) formalism is associated with the decision to express the exact energy spectrum with one of the two possible self-consistent like problems where the use of the Green operator imposes an outgoing boundary condition on the solutions of the time-independent Schrödinger equation.” Within the framework of the NHQM formalism the Green operator is described using all the discrete complex poles of the scattering matrix. That is, the NHQM Green operator, in its spectral representation, is defined as $\hat{G}(E)=\sum_{pol}\frac{|E_{pol})(E_{pol}|}{E-E_{pol}}$, were $E_{pol}$ is a complex number.
@@
Not all complex poles in NHQM are linked to resonance states; nevertheless, all resonances are indeed complex poles of the scattering matrix, which we designate as physical complex poles. These physical and non-physical discrete complex poles can serve as a basis set for calculating reaction rates for structure-less transition probabilities across a potential barrier[16]. However, in this paper, we only use the physical ones, which are devided into two categories. The occupied resonance, a transition state (TS), which corresponds to a predissociation of the activated complex, as described above. The other physical resonances decay faster than the TS resonance and result from the trapping of reactants for a relatively extended period due to vibrations perpendicular to the reaction coordinate near the saddle point in the potential energy surface (see for example Fig.5). As we will explain below, the calculations of these physical resonances can be simplified by using the nuclear adiabatic approach. This is done in order to avoid the calculations of the resonance solutions of a multi-dimensional nuclear time-independent Schrödinger equation by calculating the resonances of a modified adiabatic one-dimensional (1D) problem. Since these physical resonances are often embedded inside the 1D potential barrier (of the effective 1D adiabatic potential) we refer to them as dynamical barrier (DB) resonances. For further explanation see Ref.17.
@@
The TS predissociation resonance of the activated complex $[ABC]^{\#}$ can be coupled to one of the DB resonances with a QED mode. The TS resonance is the only one associated with the vacuum of the cavity and therefore is the only physical resonance that is occupied before tuning the distance between the two mirrors of the cavity to couple the occupied TS state with one of the selected DB resonances. Before delving into the cavity effect on chemical reactions, we must first describe the nuclear adiabatic theorem obtained from the solution of the time-independent Schrödinger equation using the reaction path Hamiltonian.
@@ III.1 The Reaction path Hamiltonian from a potential energy surface
The potential in multi-dimensional space in mass weighted coordinates is given by $V(X,Y_{1},Y_{2},...,Y_{3N-6})$, where $X$ is the reaction coordinate and $\{Y_{j}\}_{j=1,2...,3N-6}$ are perpendicular coordinates to $X$. Therefore, $(X=0,Y=0)$ is a saddle point, which is associated with an electronic transition state. Here, $N$ is the number of atoms in the $A+B$ (incoming channel), which equals to the number of atoms in $C+D$ (outgoing channel), and $63N-$ 6 is the number of internal degrees of freedom. However, only $73N-$ 7 degrees are relevant since one degree corresponds to the reaction direction. The frequency of a transition state along the reaction has an imaginary value and for the reactant/products it is zero since it corresponds to the vibration between (minimum) two species that are far apart. Therefore, approximately, we can describe this reaction path Hamiltonian (RPH) in the vicinity of the reaction coordinate where,
@@
where $\mu=m_{A}m_{BC}/(m_{A}+m_{BC})$ is the reduces mass and $V_{r.c.}(X)$, the 1D potential (corresponding to the electronic many-body solutions at different X) calculated along the reaction coordinate supports a potential barrier at $X\equiv 0$. The predissociation resonance solutions are the eigenfunctions of $\hat{H}_{RPH}$ when outgoing wave boundary conditions are imposed. The narrowest predissociation resonance with energy which is approximately as the energy in classical trajectory calculations that describe the reaction from reactants to product is associated with the transition state. Here we get into an important issue in our study. Since the TS is associated with the narrowest predissociation resonance all other resonances decay faster. The ability to couple the TS with one of the broader resonance (denoted here as DB resonance) in a dark cavity depends on the ability to have a QED mode $\hbar\omega_{cav}=Re[E_{TS}]-Re[E_{DB}]>0$. This condition is crucial for enhancement of rates of reaction in a dark cavity.
@@
For $n_{j}=0$ we get the adiabatic Hamiltonian for the ground vibrational states of the normal modes perpendicular to the reaction coordinate. In the adiabatic potential given in Eq.2 there are two type potential barriers. A static potential barriers (SB) which is associated with $V_{r.c.}(X)$ and a dynamical potential barrier (DB) which is associated with $V_{DB}(X)$. The complex poles that will be obtained in the next section are the Feshbach resonances that would be obtained from the study of the dynamics of the multi-dimensional PES as given in Eq.2. This approach has been previously used to study the collisions of hydrogen atom with $ArCL$ Van der Waals molecule[20] and for several collision reactions, such as $H+H_{2}$[21, 22]. It’s important to note that the complex physical poles are embedded inside the dynamical potential barrier that is obtained due the use of the adiabatic approximation described above, and they are not embedded in the so-called static potential barrier of $V_{r.c.}(X)$. This is a crucial point in our derivation of the conditions where the QED modes in a dark cavity are not excited (i.e., in its vacuum QED state).

### v3

# Conditions for enhancement of gas phase chemical reactions inside a dark microwave cavity

@@ Abstract
Enhancing chemical reactions, such as $A+B\to[\textit{activated complex}]^{\#}\to C+D$, in gas phase through its coupling to quantized electromagnetic modes in a dark cavity is investigated. The main result is that the enhancement of the reaction rate by a dark cavity occurs for asymmetric reactions (products different from reactants). In addition to the cavity being dark, also the reactants are required to be in their ground electronic and vibrational states, i.e., the whole system should be prepared as dark. Theoretical derivation, utilizing the non-Hermitian formalism of quantum mechanics (NHQM), provides conditions and guidelines for selecting the proper type of reactions that can be enhanced by a dark cavity. Nevertheless, the time-dependent simulations of such experiments can be carried out using the standard (Hermitian) scattering theory (while utilizing the conditions derived via NHQM). We believe that this work opens a gate to new types of studies and hopefully helps to close the gap between theory and experiments in this fascinating, relatively new field of research. As an example, we demonstrate that the asymmetric reaction rates of $O+D_{2}\to[ODD]^{\#}\to OD+D$ and $H+ArCl\to[ArHCl]^{\#}\to H+Ar+Cl$ can be enhanced by a dark cavity. On the other hand, an effect of the dark cavity on the symmetric reaction of hydrogen exchange in methane is predicted to be negligible. The applications of the theory presented here are applied to chemical reactions inside a dark microwave cavity. However, the theory is applicable to any other type of cavity and is not limited to microwave cavities only.
@@ I Motivation
The ability to control, i.e., to slow down or enhance chemical reactions, by a seemingly simple setup of reactions inside a cavity made of two parallel mirrors is fascinating. Unfortunately, currently, theory and experiment have not yet fully converged. See for example Ref.5 where they failed to reproduce the enhanced rate of cyanate ion hydrolysis reported by Hiura et al[6]. This might be due to the recent findings indicate that the cavity can introduce friction between the molecules and the solvent. This friction can enhance the reaction rate when it is low and inhibit it when it becomes significant. See for example study of the cis-trans isomerization of HONO by Sun and Vendrell[7] and references therein. In Ref. 8, Yuen-Zhou and his co-workers have shown that the ratio between rate coefficients inside and outside the cavity decreases as the energy of the lower polariton mode decreases (with increasing Rabi splitting), as expected from Marcus’s theory (see Fig. 4 in Ref. 8).
@@
The theory presented below is based on the non-Hermitian formalism of quantum mechanics (NHQM), which has previously been utilized in studying the cavity effects in chemistry. It led to the discovery of novel phenomena, such as the novel mechanism for generating high-purity single-photon emission at high repetition rates from a coherent light source[9], and the conditions enabling the upper bound polariton to penetrate into the continuum and become finite lifetime metastable state[10]. NHQM enables the calculations of the complex discrete poles of the scattering matrix, i.e., a single NHQM eigenstate instead of a Hermitian wavepacket. We will use these poles as a basis set in calculating cross sections and reaction rates. The advantage of this non-standard approach is in the ability to isolate the structures in the cross sections from the continuum background. A single time independent resonance complex pole may include all measurable information of a specific quantity in spite of the fact that in Hermitian quantum mechanics one needs to carry out wave packet dynamical calculations to describe this phenomena. This theoretical approach directs us toward identifying which kinds of collision reactions in gas phase can be enhanced by conducting them between two mirrors where the distance between them is carefully chosen. Recently we have shown that by using the complex poles as a basis set a remarkable agreement between theory and experiment was obtained. This was shown without using any fitting parameters, where the only input parameters for the cold molecular collision between triplet excited Helium and hydrogen molecule where the charges and masses of electrons and nuclei and the Planck constant[11, 12]. We expect to get similar high agreement between theory and experiment in our study of collision reactions such as $A+BC\to AB+C$ where a finite lifetime activated complex $[A...B...C]^{\#}$ is formed and its life time is associated with the complex resonance poles of the scattering matrix. Notice that these complex poles describe predissociation of metastable states of the activated complex. This is to differ from the isomerization reactions mentioned above[7] where the systems is a bound system and the poles do not appear in the complex energy plane but are on the real energy axis. This is an essential point for the theory presented below that provides the conditions for which collision reactions in gas phase (no frictions) can be enhanced in dark cavity due to the quantized field mode states in the vacuum.
@@
First, the principles of the mechanism that enable the enhancement of reactions in the gas phase by quantized radiation field modes in a cavity are described intuitively without equations using the standard (Hermitian) quantum mechanics formulation. Second, a brief overview of the fundamental concepts pertinent to this study will be provided. This will encompass the definition of discrete complex poles within the scattering matrix and its role as a basis set. In addition, it describes the predissociation-resonance metastable states of the activated complex $[ABC]^{\#}$, and the nuclear adiabatic theorem that simplified the calculations of the finite-lifetime resonances of $[ABC]^{\#}$. Second, the conditions that should be satisfied in order to enhance the rate of a reaction in a cavity will be derived. We demonstrate that the reaction rate for $O+D_{2}\to[ODD]^{\#}\to OD+D$ can be increased by placing it inside a dark cavity. The conditions for the enhancement are obtained through tuning the distances between the two cavity-mirrors. Note that cavities with different distances between the two mirrors can have multiple modes resonant with a given molecular transition (see Fig.2d in Ref.13). This tuning supports a quantized field mode that couples two resonance states, the occupied state, referred to a transition state (TS) and an unpopulated one referred to a dynamical barrier (DB) resonance state. The DB state remains unpopulated when the distance between the two mirrors is detuned. Third, we will discuss a possible feasible experiments based on the conditions derived herein. Last we conclude. In order to convince readers that it is worth the effort to explore a new theoretical concept rather than using standard approaches in the calculation of reaction rates, the results of our nonhermitian theory were introduced before delving into the formalism. This was done to attract the attention of readers to the above findings, which might help experimentalists select reactions whose rates can be enhanced inside the cavity. Notice that the applications of the theory presented here are applied to chemical reactions inside a dark microwave cavity. However, the theory is applicable to any other type of cavity and is not limited to microwave cavities only.
@@ II An intuitive (Hermitian) explanation of the cavity-induced reaction rate enhancement
Let us assume we carry out standard wavepacket calculations to describe the transition from reactants to products on a single electronic multi-dimensional potential energy surface (PES). Although in this paper we employ the non-Hermitian formalism, in which this dynamics is described differently using the non-Hermitian time-independent Hamiltonian. In this Section we do not provide a non-Hermitian explanation since the conclusions are the same. In order to cross the potential barrier that separates the reactants from the products, the collision energy must be at least equal to the energy of the barrier. This is because the transition effect due to tunneling is negligible. However, when the vibrational frequency of the colliding molecules changes substantially during the collision then the transition from reactants to products can be made at a lower collision energy. This is enabled due to transfer of energy from the translational motion along the reaction coordinate into the lowest bound state of the vibrational mode which is perpendicular to the reaction coordinate. Notice that the coupling of the molecular modes with the quantized field modes does not always enhance the coupling between the translational coordinate and the orthogonal vibrational coordinates. The just mentioned translational-vibrational coupling is enhanced only when the frequency of the bound vibrational modes varies non-monotonically at the transition state, which is located near the saddle point where a passage from reactants to products occurs. This description is equivalent to bypassing the static potential barrier along the 1D reaction coordinate but in the multidimensional potential energy surface (PES). In the multidimensional PES dynamics, the transition from reactants to products is through a saddle point, which is the transition state. When the reactants are far apart the system is described as a delocalized wavepacket, it is a continuum function along the reaction coordinate. However, around the PES saddle-point the wavepacket becomes a Gaussian like function (without a node along the coordinates, this important for identifying the transition state).
Thus, the reaction rate depends on the strength of the coupling between the translational motion along the reaction coordinate to the bound vibrational mode perpendicular to the reaction coordinate. Above we described how the strength of this coupling is due to the changes in the frequency of the vibration along the reaction coordinate. Inside the cavity the coupling of the molecular modes with the quantized field modes enhance the strength of this coupling and thereby enhance the rate of transition from reactants to products. Within NHQM it is possible to determine a suitable reaction for accelerating the reaction rate in a dark cavity without performing simulations of the experiment. In addition, it allows determination of the optimal distance between the mirrors in the cavity that will enable the optimal coupling described above.
@@ III Brief overview of the fundamental concepts pertinent to this study
Let us focus here on a collision of atom $A$ with diatom $BC$ (reactants) to form a finite lifetime activated complex $[ABC]^{\#}$, and then proceeds to yield the products $AB+C$. Reaction rates[14] and its cumulative reaction probabilities[15] can be calculated by the standard (Hermitian) quantum scattering theory. In particular, Ref.15 presents the transition from Lippmann-Schwinger formula for transition probability to Flux operator, calculated at any given divided surface. In standard quantum scattering theory the Green operator is defined as $\hat{G}^{+}(E)=\lim_{\epsilon\to 0}(E+i\epsilon-\hat{H}_{ABC})^{-1}$ where $\hat{H}_{ABC}$ is the real physical Hamiltonian for the $A+BC\to AB+C$ reaction. Notice that the term $i\epsilon$ in the Green operator introduces outgoing wave boundary condition into $|\Phi\rangle=\hat{G}^{+}(E)|\phi\rangle$ (see solution to exercise 10.1 in Ref.16). Thus, the Green operator is a non-Hermitian operator. Let us quote from Ref.17: “The branching of quantum mechanics to standard (Hermitian) formalism and non-Hermitian (NH) formalism is associated with the decision to express the exact energy spectrum with one of the two possible self-consistent like problems where the use of the Green operator imposes an outgoing boundary condition on the solutions of the time-independent Schrödinger equation.” Within the framework of the NHQM formalism the Green operator is described using all the discrete complex poles of the scattering matrix. That is, the NHQM Green operator, in its spectral representation, is defined as $\hat{G}(E)=\sum_{pol}\frac{|E_{pol})(E_{pol}|}{E-E_{pol}}$, were $E_{pol}$ is a complex number. Note that here we expand the non-Hermitian Green operator in a basis set of the complex poles of the Hamiltonian, and we should thus use the complex inner product (so-called c-product[16]) rather than the scalar product that is commonly used within the hermitian formalism. Namely, $|E_{pol})=|E_{pol}^{*}\rangle$.
@@
Not all complex poles in NHQM are linked to resonance states; nevertheless, all resonances are indeed complex poles of the scattering matrix, which we designate as physical complex poles. These physical and non-physical discrete complex poles can serve as a basis set for calculating reaction rates for structure-less transition probabilities across a potential barrier[18]. However, in this paper, we only use the physical ones, which are divided into two categories. The occupied resonance, a transition state (TS), which corresponds to a predissociation of the activated complex, as described above. The TS resonance is associated with the reaction that takes place outside of the cavity (when the two mirrors are sufficiently far apart) and can be described by a classical mechanism where the reactants have enough energy to overcome the potential barrier of the reaction. The other physical resonances decay faster than the TS resonance due to the trapping of reactants for a relatively extended time period due to bound vibrations perpendicular to the reaction coordinate near the saddle point in the potential energy surface (see for example Fig.5). Thus, even in the context of gas phase reactions, the relatively long lifetime of the activated complex arises due to crucial role played by Feshbach resonances, which result from the dynamics in the two dimensional space. To the just discussed effective potential energy barrier that results from the dynamics in 2D space we will refer later as a dynamical potential barrier (DB). As we will explain below, the calculations of these physical resonances can be simplified by using the nuclear adiabatic approach. This is done in order to avoid the calculations of the resonance solutions of a multi-dimensional nuclear time-independent Schrödinger equation by calculating the resonances of a modified adiabatic one-dimensional (1D) problem. Since these physical resonances are often embedded inside the 1D potential barrier (of the effective 1D adiabatic potential) we refer to them as dynamical barrier (DB) resonances. For further explanation see Ref.19.
@@
The TS predissociation resonance of the activated complex $[ABC]^{\#}$ can be coupled to one of the DB resonances with a quantized field mode. The TS resonance is the only one associated with the vacuum of the cavity and therefore is the only physical resonance that is occupied before tuning the distance between the two mirrors of the cavity to couple the occupied TS state with one of the selected DB resonances. Before delving into the cavity effect on chemical reactions, we must first describe the nuclear adiabatic theorem obtained from the solution of the time-independent Schrödinger equation using the reaction path Hamiltonian.
@@ III.1 The Reaction path Hamiltonian from a potential energy surface
The potential in multi-dimensional space in mass weighted coordinates is given by $V(X,Y_{1},Y_{2},...,Y_{3N-6})$, where $X$ is the reaction coordinate and $\{Y_{j}\}_{j=1,2...,3N-6}$ are perpendicular coordinates to $X$. Therefore, $(X=0,Y=0)$ is a saddle point, which is associated with a transition state as determined by electronic structure calculations. Here, $N$ is the number of atoms in the $A+B$ (incoming channel), which equals to the number of atoms in $C+D$ (outgoing channel), and $63N-$ 6 is the number of internal degrees of freedom. However, only $73N-$ 7 degrees are relevant since one degree corresponds to the reaction direction. The frequency of a transition state along the reaction has an imaginary value and for the reactant/products it is zero since it corresponds to the vibration between (minimum) two species that are far apart. Therefore, approximately, we can describe this reaction path Hamiltonian (RPH) in the vicinity of the reaction coordinate where,
@@
where $\mu=m_{A}m_{BC}/(m_{A}+m_{BC})$ is the reduces mass and $V_{r.c.}(X)$, the 1D potential (corresponding to the electronic many-body solutions at different X) calculated along the reaction coordinate supports a potential barrier at $X\equiv 0$. The predissociation resonance solutions are the eigenfunctions of $\hat{H}_{RPH}$ when outgoing wave boundary conditions are imposed. The narrowest predissociation resonance with energy which is approximately as the energy in classical trajectory calculations that describe the reaction from reactants to product is associated with the transition state. Here we get into an important issue in our study. Since the TS is associated with the narrowest predissociation resonance all other resonances decay faster. The ability to couple the TS with one of the broader resonance (denoted here as DB resonance) in a dark cavity depends on the ability to have a quantized field mode $\hbar\omega_{cav}=Re[E_{TS}]-Re[E_{DB}]>0$. This condition is crucial for enhancement of rates of reaction in a dark cavity.
@@
For $n_{j}=0$ we get the adiabatic Hamiltonian for the ground vibrational states of the normal modes perpendicular to the reaction coordinate. In the adiabatic potential given in Eq.2 there are two type potential barriers. A static potential barrier (SB), $V_{SB}(X)$, is associated with $V_{r.c.}(X)$. The dynamical potential barrier (DB), $V_{DB}(X)$, results from the bound vibrational modes that are perpendicular to the reaction coordinate $X$. The sum of these provides the effective 1D adiabatic potential. The complex poles that will be obtained in the next section are the Feshbach resonances that would be obtained from the study of the dynamics of the multi-dimensional PES as given in Eq.2. This approach has been previously used to study the collisions of hydrogen atom with $ArCl$ Van der Waals molecule[19] and for several collision reactions, such as $H+H_{2}$[22, 23]. Notice that the electronic potential surface does not depend on the mass of the nuclei, and therefore it is the same for any isotope that is used in the experiments. The use of deuterium rather of hydrogen is desirable in order to increase the number of Feshbach type predissociation resonances. It’s important to note that the complex physical poles are embedded inside the dynamical potential barrier that is obtained due the use of the adiabatic approximation described above, and they are not embedded in the so-called static potential barrier of $V_{r.c.}(X)$. This is a crucial point in our derivation of the conditions where the quantized field modes in a dark cavity are not excited (i.e., in its vacuum field state).